# When AI Evaluates Its Own Work: Validating Learner-Initiated, AI-Generated Physics Practice Problems

### v1

@@ II.3 Quality metrics
exercise-is-relevant-to-user-request Measures how well the exercise aligns with the user’s explicitly requested topic or learning objective. 2: The user’s requested topic is central.
@@
Table 2: Metrics used for both numerical and 1-out-of-NN (“radio-button”) problems.
@@
bias-free-language Language used in the exercise is neutral and respectful, free of stereotyping or discriminatory content. 1: No biases or discriminatory remarks are present.
@@
contains-harmless-extra-info The exercise includes superfluous details or narrative elements that do not confuse or contradict the solution. 1: Extra information is present but not contradictory or confusing.
@@
exercise-difficulty-level Rates how difficult the exercise is for a typical undergraduate physics student, from 0 (intuitive) to 4 (expert-level challenge). 0: Intuitive (common-sense)
@@
grounded-in-textbook The exercise aligns with the course materials. 1: Matches standard course topics.
@@
solution-not-in-problem The question refrains from providing the final answer. Hints may appear, but not the full solution path or result. 1: No final answer is revealed in the statement.
@@
asks-for-numerical-answer-only Indicates whether the task specifically demands a purely numerical answer. Answers needing symbolic expressions, proofs, or multiple-choice selections do not qualify. 1: The question clearly requests a single number (possibly with units).
@@
measurement-unit-is-clearly-stated Checks whether the exercise explicitly states the desired unit for the answer (“in meters,” “in Joules,” etc.). 1: The exercise explicitly mentions units for the final answer.
@@
distinct-misconceptions Ensures that each incorrect option in a multiple-choice question represents a different common error or misunderstanding, rather than being a rephrasing of the same mistake. 1: Each incorrect option is uniquely wrong for a specific, commonly occurring reason.
@@
single-correct-option Verifies that exactly one choice in a multiple-choice question is correct. If the question designates one right answer but all are actually incorrect, it can still be considered 1 for that design. 1: Exactly one choice is correct.
@@
Figure 4 shows an example of how these metrics were transcribed within our LLM-as-a-judge pipeline as a JSON structure. Together with the exercise itself, these instructors are fed into a prompt template that expects a structured JSON output, which is then processed by the evaluation pipeline.
@@
II.4 Ground truth
To establish the ground truth, the questions and dialogues were evaluated on the quality metrics by one of the authors (G.K.), who taught introductory physics courses for over two decades.
II.4.1 Skewedness
Although we collected a total of $N=543$ generated exercises, the class distribution for some metrics was highly skewed — indeed, certain classes $C_{i}$ were underrepresented or even absent. To ensure that per-class estimates (e.g., precision or recall for each $C_{i}$) have sufficiently low sampling variability before treating them as definitive, we adopted the conventional rule of requiring at least 30 observations per class. This “30-sample rule” is a rule of thumb based on the central limit theorem: binomial proportions begin to approximate the normal distribution when $n\gtrsim 30$, yielding 95 % confidence interval half-widths on the order of $\pm 18$ percentage points or less [58, 59].
@@ II.4.1 Skewedness
In all 543 exercises, there was no case of biased language or offensive content. Gender-neutral language was used in all cases. Subjectively, most problems were context-independent, focusing on circuit elements or blocks on slopes; these theoretical, “clean room” situations were not at risk of biases.
@@
There were only five cases of exercises that were not grounded in the textbook, and those turned out to be hallucinations like a biology problem about mitosis or a problem about the seasons on Earth (“hallucinations” in this context refers to instances where an LLM generates plausible-sounding but inaccurate or entirely fabricated information [60]).
@@
There was only one exercise that had unrealistic data, namely a problem with a freight train that has a total mass of 5,000 kg, i.e., five tons, which would be the mass of a small truck.
@@
In only seven cases was the generated exercise not relevant to the user request, and among those were the five cases that were not grounded in the textbook.
II.4.2 Interrelatedness
As Fig. 5 shows, several metrics are correlated with each other [61]. This may allow us to later use a metric that is more reliable as a proxy for relevant, but less reliable, metrics. Added here is the variable chosen if the student chose to work on this problem in the forced-choice shown in Fig. 1.
Figure 5: Correlations between the metrics in a force-directed Fruchterman-Reingold representation [62]. Positive correlations are shown in green, negative correlations in red. The thickness of the lines indicates the Spearman correlation strength [61]. Any edges corresponding to an absolute correlation value <0.2<0.2 and isolated nodes have been removed. All shown correlations are highly significant (p <0.001<0.001). The clusters identified in Sect. II.4.2 are highlighted in green, blue, and yellow, respectively; the variables chosen and problem-type-is-radio are also connected to Cluster A, but not determined by the exercise-validation step in Fig. 3.
Three highly significant ($p<0.001$) clusters emerge from the force-directed graph in Fig. 5:
@@ II.4.2 Interrelatedness
A salient negative hub is problem-type-is-radio, which correlates inversely with nearly every node in Cluster A, most pronounced with includes-solution-strategy ($\rho=-0.71$) and Bloom level ($\rho=-0.68$). This mirrors the feature-importance result that radio items are less likely to be selected when a more structured, quantitatively framed alternative is available.
II.5 Considered LLMs
@@
II.6 Measures of metrics judgment performance
@@
When comparing the judged values of the metrics to the ground truth, we took into account the unbalanced nature of our dataset by considering macro-level scores:
@@
II.7 Measures of user choices
In the user study, when students requested an exercise from the chatbot, they were presented with two alternative problems to choose from (see Figure 1); this choice is captured by the metric chosen. In addition to Spearman correlations [61] of metrics to chosen, we also analyzed which metrics predict chosen. We employed random forest models (a type of ensemble machine learning algorithm) [68], which we trained for each exercise type (11-out-of-NN (radio-button) and numerical exercises).
@@ II.7 Measures of user choices
II.8 Use of AI
@@ III.1 Performance of LLMs
Given that our dataset contains unbalanced metrics, we evaluated macro-level quality measures when comparing the LLM-judgments to the ground truth. Table 5 shows the results.
Table 5: Macro-level quality indicators of the models in Sect. II.5 with the highest F1-score, sorted by F1-score. The superscripts “num” and “radio” indicate metrics only available for numerical and radio-button problems, respectively. The superscript “multi” indices metrics with more than two classes. Finally, the asterisk indicates confirmatory metrics, i.e., fulfilling the minimum-sample rule in Sect. II.4. The metrics above the upper horizontal line (up to grounded-in-textbook) have artificially high F1-values due to lacking variance in the sample. Metrics below the lower horizontal line have F1<0.6<0.6, which is generally considered poor agreement between model and ground truth.
@@
Across our suite of evaluation metrics, some scores appear deceptively perfect. For instance, both bias-free-language and realistic-data-used achieved an F1 of 1.00. However, these ceiling values arise because the positive class in each case is virtually non-existent: the models simply default to the prevailing (negative) label, and in the rare instance of a mislabel — just one incorrect data usage case — our system caught it flawlessly. By contrast, only three checks, includes-solution-strategy, measurement-unit-is-clearly-stated, and llm-solution-is-correct, surpass our $0.80$ reliability threshold, each scoring in the mid-$0.80$s and satisfying our minimum-sample criterion.
Meanwhile, the core problem-statement metrics show moderate utility but leave substantial room for improvement. Metrics such as task-is-solvable, task-is-clear, and task-is-specific-and-complete cluster around F1 values near 0.570.610.57-0.61, meaning they correctly flag most instances but still miss roughly 40 % of edge cases. Multi-category labels perform even more poorly: all five categorical rubrics, including bloom-level-of-exercise at $0.40$ and exercise-difficulty-level at $0.23$, fall well below the $0.50$-mark, underscoring that human oversight remains indispensable whenever nuanced distinctions are required. Notably, the exercise-is-relevant-to-user-request metric scores just $0.47$ overall: while the models excel at identifying irrelevant exercises, they struggle to distinguish between partial and perfect relevance. Simplifying this to a binary classification might streamline future evaluations without sacrificing the robust detection of off-topic items.
When we aggregate across F1 scores to see which model “wins” most often, the o3-mini/low variant leads with $11$ victories (44 %), followed by GPT-4o with 9 wins (36 %) and GPT-4o-mini with 5 wins (20 %). Although o3-mini/low holds the top spot, more than half of the winning metrics still favor GPT-4o-based models. Cost considerations further complicate the picture: at the time of writing (Spring 2025), o3-mini/low runs roughly seven times more expensive than GPT-4o-mini but remains cheaper than full GPT-4o, striking a compelling balance of performance and price.
@@ III.2 Relevance of metrics to user choices
Tables 6 and 7 show the outcomes of our analysis as to which metrics influence user choices when presented with two problem alternatives as shown in Fig. 1. For both models, the metrics expected-bloom-level and expected-student-difficulty-level were dropped, as in more than 30% of the cases, their ground truth could not be discerned from the prompt and its context.
@@
Table 6: Random forest feature importances for predicting whether any given 11-out-of-NN (“radio-button”) exercise was chosen by the student when presented with two exercises. The asterisk indicates confirmatory metrics, i.e., fulfilling the minimum-sample rule in Sect. II.4. The horizontal line indicates a threshold of $0.1$.
@@
Table 7: Random forest feature importances for predicting whether any given numerical exercise was chosen by the student when presented with two exercises. The asterisk indicates confirmatory metrics, i.e., fulfilling the minimum-sample rule in Sect. II.4. The horizontal line indicates a threshold of $0.1$.
@@
Finally, Table 8 shows the random forest importances of metrics that are common to both exercise types, adding an additional metric problem-type-is-radio to distinguish the two. The model was trained with 372 training and 93 test samples. It reaches a test accuracy of 73.1 % (CV $0.656\pm 0.029$, OOB 66.7 %). Its importance ranking is headed by includes-solution-strategy (0.301) and the format flag (0.192); conceptual-depth metrics (bloom-level-of-exercise, 0.130; exercise-difficulty-level, 0.115) follow next, with llm-solution-is-correct (0.098) completing the top five. As it turns out, problem-type-is-radio is the only metric strongly influencing user choice that is negative: students prefer numerical problems.
Table 8: Random forest importances of metrics applicable to both radio and numerical exercises. The asterisk indicates confirmatory metrics, i.e., fulfilling the minimum-sample rule in Sect. II.4. The horizontal line indicates a threshold of $0.1$.
@@
(includes-solution-strategy) ranks first or second in every model.
@@
Both Bloom level and difficulty level enter the importance lists, signaling that students notice cues related to the intellectual demand of an exercise.
The importance of llm-solution-is-correct, which barely did not make the cut in all three scenarios, is initially puzzling, as the student would not know this before having chosen and attempted the problem. However, this metric is strongly correlated to the measures of surface structure and clarity, namely the “Cluster B — Task Clarity & Feasibility” discussed in Sect. II.4.2; as also Fig. 5 shows, the metric thus gets elevated alongside these immediately visible metrics.

### v2

@@ II.4 Validity metrics
exercise-is-relevant-to-user-request Measures how well the problem aligns with the user’s explicitly requested topic or learning objective. 2: The user’s requested topic is central.
@@
Table 2: Metrics used for both numerical and 1-out-of-NN (“radio-button”) problems. In our technical implementation, the terms exercise, task, and (practice) problem were used interchangeably.
@@
bias-free-language Language used in the problem is neutral and respectful, free of stereotyping or discriminatory content. 1: No biases or discriminatory remarks are present.
@@
contains-harmless-extra-info The problem includes superfluous details or narrative elements that do not confuse or contradict the solution. 1: Extra information is present but not contradictory or confusing.
@@
exercise-difficulty-level Rates how difficult the problem is for a typical undergraduate physics student, from 0 (intuitive) to 4 (expert-level challenge). 0: Intuitive (common-sense)
@@
grounded-in-textbook The problem aligns with the course materials. 1: Matches standard course topics.
@@
solution-not-in-problem The problem refrains from providing the final answer. Hints may appear, but not the full solution path or result. 1: No final answer is revealed in the statement.
@@
asks-for-numerical-answer-only Indicates whether the task specifically demands a purely numerical answer. Answers needing symbolic expressions, proofs, or multiple-choice selections do not qualify. 1: The problem clearly requests a single number (possibly with units).
@@
measurement-unit-is-clearly-stated Checks whether the problem explicitly states the desired unit for the answer (“in meters,” “in Joules,” etc.). 1: The problem explicitly mentions units for the final answer.
@@
distinct-misconceptions Ensures that each incorrect option in a multiple-choice problem represents a different common error or misunderstanding, rather than being a rephrasing of the same mistake. 1: Each incorrect option is uniquely wrong for a specific, commonly occurring reason.
@@
single-correct-option Verifies that exactly one choice in a multiple-choice problem is correct. If the problem designates one right answer but all are actually incorrect, it can still be considered 1 for that design. 1: Exactly one choice is correct.
@@
Figure 5 shows an example of how these metrics were transcribed within our LLM-as-a-judge pipeline as a JSON structure. This description of the respective evaluation metric, the problem being evaluated, and the evaluation instructions are input into a prompt template to elicit a structured JSON response from the LLM, which the evaluation pipeline then processes.
@@
II.5 Ground truth
To establish the ground truth, the problems and dialogues were evaluated on the quality metrics by one of the authors (G.K.; T.G. is a computer scientist), who taught introductory physics courses for over two decades. This manual rating of $\approx 10,000$ metrics took several days.
II.5.1 Class imbalance (“skewedness”)
A class $C_{i}$ denotes one of the discrete, mutually exclusive outcome labels that a given metric can take. For example, for task-is-solvable the classes are $\{0,1\}$; for bloom-level-of-exercise the classes are $\{1,2,3,4\}$; and for exercise-is-relevant-to-user-request the classes are $\{0,1,2\}$.
Although we collected a total of $N=543$ generated problems, the human-rated class distribution for some metrics was highly skewed — indeed, certain classes $C_{i}$ were underrepresented or even absent. To ensure that per-class estimates (e.g., precision or recall for each $C_{i}$) have sufficiently low sampling variability before treating them as definitive, we adopted the conventional rule of requiring at least 30 observations per class. This “30-sample rule” is a rule of thumb based on the central limit theorem: binomial proportions begin to approximate the normal distribution when $n\gtrsim 30$, yielding 95 % confidence interval half-widths on the order of $\pm 18$ percentage points or less [66, 67].
@@ II.5.1 Class imbalance (“skewedness”)
In all 543 problems, there was no case of biased language or offensive content. Gender-neutral language was used in all cases. Subjectively, most problems were context-independent, focusing on circuit elements or blocks on slopes; these theoretical, “clean room” situations were not at risk of biases.
@@
There were only five cases of problems that were not grounded in the textbook, and those turned out to be hallucinations like a biology problem about mitosis or a problem about the seasons on Earth (“hallucinations” in this context refers to instances where an LLM generates plausible-sounding but inaccurate or entirely fabricated information [68]).
@@
There was only one problem that had unrealistic data, namely a problem with a freight train that has a total mass of 5,000 kg, i.e., five tons, which would be the mass of a small truck.
@@
In only seven cases was the generated problem not relevant to the user request, and among those were the five cases that were not grounded in the textbook.
II.5.2 Interrelatedness
We now investigate how metrics in the ground truth are correlated with each other. Added here is the variable chosen if the student chose to work on this problem in the forced-choice shown in Fig. 2; this metric reflects student preferences and the problem’s perceived alignment with desired features (one might argue that students may not know “what’s good for them,” however, we are modeling a scenario where students are motivated by studying for an exam, not a homework scenario where they might choose the path of least resistance). This may allow us to later use a metric that is more reliable as a proxy for relevant, but less reliable, metrics.
The force-directed Fruchterman–Reingold representation [69] in Fig. 6 places the metrics, represented as nodes, in two dimensions by minimizing the equivalent of potential energy where adjacent nodes attract (spring forces proportional to their pairwise Spearman correlations $\varrho$ [70]), and all nodes mutually repel (see Ref. [71] for a physics derivation of this layout); the result is a geometry in which strongly connected variables appear closer together and clustering becomes visually apparent.
Figure 6: Correlations between the metrics in a force-directed Fruchterman-Reingold representation [69]. Positive correlations are shown in green, negative correlations in red. The thickness of the lines indicates the Spearman correlation strength [70]. Any edges corresponding to an absolute correlation value $\varrho<0.2$ and isolated nodes have been removed. All shown correlations are highly significant (p <0.001<0.001). The clusters identified in Sect. II.5.2 are highlighted in green, blue, and yellow, respectively; the variables chosen and problem-type-is-radio are also connected to Cluster A, but not determined by the problem-validation step in Fig. 4.
Three clusters emerge from Fig. 6, based on pairwise highly significant Spearman correlations $\varrho\geq 0.2$ (p <0.001<0.001) :
@@ II.5.2 Interrelatedness
A salient node that ties together other nodes (a “hub”) is problem-type-is-radio, which correlates inversely with nearly every node in Cluster A (“negative hub”), most pronounced with includes-solution-strategy ($\rho=-0.71$) and Bloom level ($\rho=-0.68$). This mirrors the feature-importance result that radio items are less likely to be selected when a more structured, quantitatively framed alternative is available.
II.5.3 Subjective observations
While establishing the ground truth, all problems generated in this study had to be read and solved. Subjectively, they contained a disproportionate number of run-of-the-mill standard problems, which were generally solvable in one or two steps. The problems generally lacked sophistication or interest, similar to typical end-of-chapter problems in non-research-based textbooks. The scenarios and tasks were at times repetitive with only slight variations, and they were generally not situated in authentic contexts.
II.6 Considered LLMs
@@
Because the study ran on the shared Ethel platform, whose resources simultaneously served 3000 students across 20 courses at that time, we were unable to separate its actual computational cost from general operations; we are thus unable to assign a price-tag to the mechanisms under investigation.
II.7 Measures of metrics judgment performance
@@
When comparing the judged values of the metrics to the ground truth, we took into account the unbalanced nature of our dataset by considering macro-level scores, which are the arithmetic mean of the individual class scores:
@@
II.8 Measures of user choices
In the user study, when students requested a problem from the chatbot, they were presented with two alternative problems to choose from (see Figure 2); this choice is captured by the metric chosen. In addition to Spearman correlations [70] of metrics to chosen, we also analyzed which metrics predict chosen. We employed random forest models (a type of ensemble machine learning algorithm) [77], which we trained for each problem type (11-out-of-NN (radio-button) and numerical problems).
@@ II.8 Measures of user choices
II.9 Thresholds
We are setting the following thresholds for our three quality goals:
Reliable
For items that satisfy our minimum-sample rule (Sect. II.5), we quantify reliability by macro-averaged precision, recall, F1, and accuracy computed per class and then averaged across classes. We treat F1 $\geq 0.80$ as strong agreement (“reliable”), $0.60\leq$ F1 <0.80<0.80 as moderate, and F1 <0.60<0.60 as poor; metrics that violate the per-class $n\!\geq\!30$ rule are labeled exploratory regardless of score (Sect. II.5).
Relevant
Relevance is established by (i) a statistically significant association with chosen (Spearman $|\rho|\!\geq\!0.20$, p<0.001p\!<\!0.001; Sect. II.5.2) and/or (ii) non-trivial predictive contribution in held-out random-forest models (feature importance $\geq 0.10$), with converging qualitative support from exit-survey comments. In other words, relevant metrics either predict or robustly correlate with the options learners pick.
Automatically assessable
Practically, we restrict to metrics that can be scored with one (or a small fixed number of) commodity LLM calls per item, keeping latency and cost compatible with a conversational interface, and that succeed robustly in structured output parsing across the dataset.
II.10 Use of AI
@@ III.1 Performance of LLMs
Table 5 shows the macro-level quality measures when comparing the LLM-judgments to the ground truth.
Table 5: Macro-level quality indicators of the models in Sect. II.6 with the highest F1-score, sorted by F1-score. The superscripts “num” and “radio” indicate metrics only available for numerical and radio-button problems, respectively. The superscript “multi” indices metrics with more than two classes. Finally, the asterisk indicates confirmatory metrics, i.e., fulfilling the minimum-sample rule in Sect. II.5. The metrics above the upper horizontal line (up to grounded-in-textbook) have artificially high F1-values due to lacking variance in the sample. Metrics below the lower horizontal line have F1<0.6<0.6, which is generally considered poor agreement between model and ground truth.
@@
Across our suite of evaluation metrics, some scores appear deceptively perfect. For instance, both exploratory metrics bias-free-language and realistic-data-used (see Subsec. II.5.1) achieved an F1 of 1.00. However, these ceiling values arise because the positive class in each case is virtually non-existent: the models simply default to the prevailing (negative) label, and in the rare instance of a mislabel — just one incorrect data usage case — our system caught it flawlessly. By contrast, only three checks, includes-solution-strategy, measurement-unit-is-clearly-stated, and llm-solution-is-correct, surpass our $0.80$ reliability threshold, each scoring in the mid-$0.80$s and satisfying our minimum-sample criterion.
In this context, it is important to emphasize that llm-solution-is-correct is not just “AI works because AI says it works.” In our ground truth, the metric was determined by the human expert. In our current implementation, it aligns well but not perfectly with independent LLM judgements of the LLM-generated problems: the problem was first generated and then subsequently and independently judged with results comparable to human judgement.
Meanwhile, the core problem-statement metrics show moderate utility but leave substantial room for improvement. Metrics such as task-is-solvable, task-is-clear, and task-is-specific-and-complete cluster around F1 values near 0.570.610.57-0.61, meaning they correctly flag most instances but still miss roughly 40 % of edge cases. Multi-category labels perform even more poorly: all five categorical metrics, including bloom-level-of-exercise at $0.40$ and exercise-difficulty-level at $0.23$, fall well below the $0.50$-mark. Figure 7 shows the distributions of expected-student-difficulty-level and exercise-difficulty-level in both the expert-rating and the AI-judgement to illustrate these discrepancies in these two worst-performing metrics. In the majority of cases, the human expert was unable to discern the expected difficulty from the prompt and context, while the AI deemed most prompts and contexts to expect a straightforward problem; while the option of ‘N/A” was available, the AI defaulted to “foundational.”
Figure 7: Distributions of expected-student-difficulty-level and exercise-difficulty-level as judged by the human expert and the AI (using o3-mini/low as example); see Table 1 (human) and Table 2 (AI) for the definitions of the difficulty levels.
Notably, the exercise-is-relevant-to-user-request exploratory metric scores just $0.47$ overall: while the models excel at identifying irrelevant problems, they struggle to distinguish between partial and perfect relevance. Simplifying this to a binary classification might streamline future evaluations without sacrificing the robust detection of off-topic items.
When we aggregate across scores to see which model “wins” most often in terms of having the highest macro F1 score, the o3-mini/low variant leads with $11$ victories (44 %), followed by GPT-4o with $9$ wins (36 %) and GPT-4o-mini with $5$ wins (20 %). Although o3-mini/low holds the top spot, more than half of the winning metrics still favor GPT-4o-based models — the reasoning model does not significantly outperform the non-reasoning one. Cost considerations further complicate the picture: at the time of writing (Spring 2025), o3-mini/low runs roughly seven times more expensive than GPT-4o-mini but remains cheaper than full GPT-4o, striking a compelling balance of performance and price.
@@ III.2 Relevance of metrics to user choices
Tables 6 and 7 show the outcomes of our analysis as to which metrics influence user choices when presented with two problem alternatives as shown in Fig. 2. This user choice is not a measure of problem quality, as the students would likely not yet have attempted to solve the problem at the time they made the choice, and as they may not have enough mastery of the content yet to judge its appropriateness; instead, it reflects students’ preferences and the problem’s perceived alignment with, in their eyes, desirable properties. For both models, the metrics expected-bloom-level and expected-student-difficulty-level were dropped, as in more than 30% of the cases, their ground truth could not be discerned by the human rater from the prompt and its context; for example, a prompt like “please give me a problem involving Kirchhoff rules” provides no information about the desired Bloom or difficulty levels, and thus the associated metrics were coded as missing or non-applicable (“N/A”) by the rater.
@@
Table 6: Random forest feature importances for predicting whether any given 11-out-of-NN (“radio-button”) problem was chosen by the student when presented with two problems. The asterisk indicates confirmatory metrics, i.e., fulfilling the minimum-sample rule in Sect. II.5. The horizontal line indicates a threshold of $0.1$.
@@
Table 7: Random forest feature importances for predicting whether any given numerical problem was chosen by the student when presented with two problems. The asterisk indicates confirmatory metrics, i.e., fulfilling the minimum-sample rule in Sect. II.5. The horizontal line indicates a threshold of $0.1$.
@@
Finally, Table 8 shows the random forest importances of metrics that are common to both problem types, adding an additional metric problem-type-is-radio to distinguish the two. The model was trained with 372 training and 93 test samples. It reaches a test accuracy of 73.1 % (CV $0.656\pm 0.029$, OOB 66.7 %). Its importance ranking is headed by includes-solution-strategy (0.301) and the format flag (0.192); conceptual-depth metrics (bloom-level-of-exercise, 0.130; exercise-difficulty-level, 0.115) follow next, with llm-solution-is-correct (0.098) completing the top five. As it turns out, problem-type-is-radio is the only metric strongly influencing user choice that is negative: students prefer numerical problems.
Table 8: Random forest importances of metrics applicable to both radio and numerical problems. The asterisk indicates confirmatory metrics, i.e., fulfilling the minimum-sample rule in Sect. II.5. The horizontal line indicates a threshold of $0.1$.
@@
(includes-solution-strategy) ranks highly in every model and is the leading predictor when considering both problem types.
@@
Both Bloom level and difficulty level enter the importance lists, signaling that students notice cues related to the intellectual demand of a problem.
The importance of llm-solution-is-correct, which barely did not make the cut in all three scenarios, is initially puzzling, as the student would not know this before having chosen and attempted the problem (except in cases where they may have deemed a problem obviously “unsolvable” because of apparent internal inconsistencies or apparently no correct answer options). In fact, one might argue that this metric should not have been included in this particular analysis, as it is generally latent in this context. However, this metric is strongly correlated to the measures of surface structure and clarity, namely the “Cluster B — Task Clarity & Feasibility” discussed in Sect. II.5.2 as confounding factors; as also Fig. 6 shows, the metric thus gets elevated alongside these immediately visible metrics. We are not claiming causation, but in our context, correlation is sufficient to mark it as a useful factor.